# Can large language models generalize analogy solving like children can?

### v2

@@ 3 Results
We use mixed ANOVAs to (1) compare performance between our between-subjects participant groups (Adults, Children, and each of the LLMs) on the Latin alphabet and (2) test whether each participant group could generalize analogy solving by performing similarly across alphabets (i.e., our repeated within-subjects factor).
@@
We expected LLMs to be able to solve letter-string analogies with the Latin alphabet at the same level as adults (Webb et al.,, 2023) and that both adults and LLMs would outperform children (Thibaut and French,, 2016) (hypotheses H1a-c). Similar to what we expected, adults and some LLMs, except Google’s Gemma-2 27B and Anthropic’s Claude 3.5, performed better than children in the Latin alphabet domain. Open AI’s GPT-4o performed similarly to adults, followed closely by Meta’s Llama-3.1 405B. See Figure 3 and Tables 2 and 3 for more detailed results.
Table 2:
Descriptive statistics on letter-string analogy performance by Participant Group and Alphabet
@@
Table 3:
Post hoc t-test results: Participant group performance comparison with children in Latin domain
Alphabet Group 1 Group 2 n1 n2 t p Adjusted
Latin Children Adults 42 62 -6.57 <.001
Latin Children Claude-3.5 42 54 -1.35 1.0
Latin Children Gemma-2 27B 42 54 0.471 1.0
Latin Children GPT-4o 42 54 -5.41 <.001
Latin Children Llama-3.1 405B 42 54 -3.85 =.001
3.2 RQ2: How well do adults, children and LLMs generalize letter-string analogy solving from the Latin (baseline) domain to the Greek (near) and Symbol (far) domains?
We expected adults and children to generalize analogy solving to other domains and therefore perform similarly across domains (Doumas et al.,, 2022) (hypotheses H2a-b). The experiment was designed to test whether LLMs could also generalize to other domains. Given the mixed results in previous research we had no clear evidence to predict how LLMs would perform on the near (Greek) and far (Symbol) letter-string analogy domains, but we suspected that LLM performance would degrade in less familiar domains.
Indeed our results indicate that adults and children perform similarly across alphabets (see Figure 3). But, for the four LLMs we tested, Anthropic’s Claude-3.5, Google’s Gemma-2 27B, Open AI’s GPT-4o, and Meta’s Llama-3.1 405B, performance indeed degraded in less familiar alphabets (ANOVA results shown in Table 4). More specifically, for each model, performance degraded significantly from the Latin to Greek alphabet (posthoc Bonferonni-corrected t-test results all p<.001, except for Llama-3.1 405Bp = 0.012) and then again from the Greek alphabet to the Symbol list (posthoc Bonferonni-corrected t-test results all p<.001).
Table 4:
Post hoc ANOVA Results for main Alphabet effect on letter-string performance by Participant Group
@@
Adults alphabet 1.59 96.9 0.95 1.000
Children alphabet 2.00 76.0 0.27 1.000
Claude-3.5 alphabet 1.65 87.6 29.5 <.001
Gemma-2 27B alphabet 2.00 106.0 88.2 <.001
GPT-4o alphabet 2.00 100.0 55.0 <.001
Llama-3.1 405B alphabet 1.70 90.1 135.0 <.001
Figure 3:
Performance comparison of human and LLM participants on letter-string analogies. Mean proportion correct scores are shown for each group across Latin (baseline), Greek (near transfer), and Symbol (far transfer) domains. Error bars represent standard error of the mean. Both adult and child participants maintained consistent performance across domains, while LLMs showed progressive performance degradation from baseline to far transfer conditions.
@@ 3.3 RQ3: Why can’t LLMs generalize letter-string analogy solving like children?
3.4 Performance by Item
To understand why the LLM’s had trouble generalizing letter-string analogy solving to the Greek and Symbol domains we first look at their performance per item as this may give insight into which rules were easier and more difficult for the LLMs to apply. Table 5 shows an overview. Here we see that the LLMs and humans perform best on item 1, that involves only the successor transformation, and worst on item 5, that involves both the predecessor transformation and repetition generalization. Because, item 2, also involves the same repetition rule as item 5, but was solved better by LLMs and children, it seems like the predecessor rule is what gives both LLMs and children the most trouble. The other item people and LLMs have relatively more trouble with is item 3. This item involves the second successor rule. In sum, the predecessor and second successor rules appear to be the most difficult rules from our item set for people and LLMs to apply.
Table 5:
Mean proportion correct (SD) by Participant Group for each Item
@@ 3.4 Performance by Item
3.5 Next-Previous Letter Task
To investigate this further, we designed the Next-Previous Letter Task to check that the LLMs had the requisite domain knowledge of predecessor, successor and second successor to perform our letter-string analogy task. The task involved providing an ordered list of letters/symbols and asking the LLMs what the previous and next letters were given a specific letter. We did this 5 times using an optimized prompt requesting to identify the letter: one before, two before (not required in our item set), one after and two after, resulting in 20 items total. The exact prompts and items can be found in the Appendix D.
As can be seen in Figure 4, all models do best when asked to identify the next or previous letter and worse when it concerns identifying the second successor and second predecessor. Furthermore, Claude-3.5 performed well and similarly in all three domains, which is in contrast to its letter-string analogy performance that degrades from baseline to near to far domains. Similarly, GPT-4o performs well in the Latin and Greek domain, but in the letter-string task, its performance degrades from Latin to Greek. For Llama-3.1 405B transfer from the Latin to Greek to Symbol domain is similar across tasks, where in both tasks it does well with the Latin and Greek alphabets, but not the Symbol alphabet. Gemma-2 27B’s performance is surprisingly more spotty here in the Greek domain than the Symbol domain. So, these results could explain why the LLMs have trouble with item 3, involving the second successor, but the results do not explain why they have trouble with item 5 involving the predecessor.
Figure 4:
LLM performance by rule type across alphabet domains. All LLMs successfully apply rules in the Latin alphabet and maintain performance on repetition rules across domains. Performance degrades significantly for predecessor and second successor rules in the Symbol domain. Error bars represent standard error of the mean.
@@ 3.5.1 Rule Check Task
To better pinpoint why the LLMs had difficulty generalizing to the Greek and Symbol alphabet domains, we created a simplified item set that explicitly tested each rule used in the human item set in isolation. The rules were: (1) successor_1, the next letter; (2) successor_2, letter two places after; (3) predecessor_1, the previous letter; (4) predecessor_2, letter two places before; (5) repetition_1, repeating the last letter and (6) repetition_2, repeating both letters. Each rule was tested five times. Example items for the Latin alphabet are shown in Table 6. LLM system and instruction prompts were exactly the same as with the original items of our letter-string task.
Table 6:
Rule Check Task: Example Items From the Latin Alphabet.
@@ A B C D Rule AB
As can be see in Figure 5, the LLMs we tested can solve all rules in the Latin alphabet and have no problem with repetition rules in the Greek and Symbol domains. The successor and predecessor rules were solved to differing degrees in the Greek alphabet, with Claude-3.5 performing best followed by GPT-4o. All models had trouble with the successor_2 and predecessor rules in the Symbol alphabet, where only the successor_1 rule sometimes formed an exception. This makes sense given the predict-the-next-token goal that LLMs are trained on (McCoy et al.,, 2024).
Figure 5:
Rule-specific performance across alphabet domains for LLM participants. Mean accuracy scores are shown for each rule type (successor, predecessor, and repetition) across Latin (baseline), Greek (near transfer), and Symbol (far transfer) domains. Error bars represent standard error of the mean. LLMs maintain near perfect performance on all rules in Latin and on repetition rules across domains, but show significant degradation on predecessor and second successor rules in the Symbol domain.
@@ 3.5.2 Error Analysis
In general, when solving letter-string analogies there are often multiple rules that could underlie the change from A to B Hofstadter and Mitchell, (1994). In the very short strings that we use there is less ambiguity about the rule, than in longer strings. In our case, there are generally only two clearly correct responses. We considered the rules that people would generally prefer when responding, to be “correct”, such as if ab changes to ac, then gh changes to gi. However, the literal rule of replacing the last letter with c, with response gc could also be considered correct.
@@
To examine errors in more detail we created a set of categories based on those from (Lewis and Mitchell,, 2024) and extended these to account for common errors in children Stevenson and Hickendorff, (2018). In the Literal rule category, the change from A to B is literally copied to C such as a b : a c c :: g h : g c c rather than providing the more common response of g i i. In the One rule category, the response is partially correct, but only (part of) one of the rules in the problem was applied, such as in responses to the previous example, g h h (only repetition applied) or g i (only successor applied). Partially correct responses are common in children when problem load supersedes processing capacity Stevenson and Hickendorff, (2018). In the Incorrect rule category, one of the other rules from our item set (i.e., successor, predecessor, repetition) was applied; for example, if the successor rule was used instead of the predecessor rule. For the Copy rule, the A, B or C term was copied as copying the C-term is common in young children (Stevenson and Hickendorff,, 2018; Opiełka et al.,, 2024). Finally, all remaining erroneous responses were placed in the Other rule category. Given that our task was less complex than in (Lewis and Mitchell,, 2024) (i.e., shorter strings, fewer rules), we were able to automatically code these categories.
Table 7:
Proportions of error categories by participant group
@@
Note. 5% of children’s responses were empty and are not included in the proportions shown.
What we see is that adults and children did not use the Literal rule, whereas all models used it sometimes, and for Gemma-2 27B and GPT-4o the Literal rule was one of the most common error types. The One rule was used most often in errors by Claude-3.5. The Incorrect and Copy rules were not used very often by people or models. And the Other rules were used most often by all, except Claude-3.5. Figure 6 shows a break down of Other rule use by Alphabet. Adults use an Other rule most often across the board, but they also have the fewest errors. Children’s Other errors are also high across alphabets. Interestingly, GPT-4o’s Other responses are generally found in the Latin alphabet, while those for the other three LLMs are most prevalent in the Symbol alphabet.
Figure 6:
Analysis of erroneous responses by participant group and alphabet domain. Mean frequency of Other rule application is shown for children, adults, and LLMs across Latin, Greek, and Symbol alphabets. Error bars represent standard error of the mean.
@@
For each response we computed the Levenshtein string distance, also known as optimal string alignment distance, from the expected “correct” response to the given response. This distance counts the minimum number of edit operations (insertion, deletion, substitution) needed to change one string into the other. Here we investigate whether there are differences in mean Levenshtein distance between adults, children and LLMs for “erroneous” responses. Figure 7 shows that that the Levenshtein distance for “erroneous” responses is greater for children on all alphabets than for LLMs. For adults, this is only the case for the Symbol alphabet. For, LLMs the Levenshtein distance hovers just under the 2 for all alphabets. Note also that the standard errors for LLMs are also much smaller, but this is because the adults and sometimes children (Greek, Symbol alphabets) had far fewer “erroneous” responses to sample from. These results tells us that when children provide “erroneous” answers their responses tended to differ largely from the expected response. For example, three children responded ‘m m‘ to the item ‘If c d changes to b d, what does m m n n change to?’, which has a Levenshtein distance of 6 from the expected response ‘l l n n‘. The LLMs tended to provide 1 or 2 expected letters and 1 or 2 unexpected ones. For example, on the same item (and its variants) six GPT-4o runs provided ‘l m m n‘ as a response, with a Levenshtein distance of 2 from the expected response.
Figure 7:
String distance comparison of incorrect responses across participant groups and domains. Mean string distance between incorrect and expected responses is shown for children, adults, and LLMs across Latin, Greek, and Symbol alphabets. Error bars represent standard error of the mean. Children’s erroneous responses tend to differ more from the expected response than those of LLMs for all alphabets. Adults erroneous responses only differ more for the Symbol alphabet.
@@ 4 Discussion
Our main finding is that the LLMs we tested, using similar prompts given to children, were not able to generalize letter-string analogy solving like children can. LLMs perform at or above the level of children on letter-string analogies in the familiar Latin alphabet, but their performance on these same problems reduces somewhat when using the Greek alphabet (near transfer) and deteriorates almost entirely when using our Symbol alphabet (far transfer).
@@
Why can’t LLMs generalize when solving letter-string analogies? For some models, this appears to be because they were unable to meet underlying requisites, such as indicating the predecessor or second successor. This would make sense given the predict-the-next-token goal that LLMs are trained on McCoy et al., (2024). We tested this using the Next-Previous letter task, where models were given an ordered list of letters or symbols and asked to identify the (second) successor or predecessor to a given letter or symbol. These results could explain why the LLMs have trouble with the second successor, however the models had little trouble identifying the predecessor in this task. So, these results do not fully explain why LLM analogy solving performance degrades from the Latin to Greek to Symbol domains.
@@
The problem with LLM’s transfer from the Latin to other domains seems to lie in that the alphabet is too “unfamiliar”. What we mean here is that the conceptual abstraction of what constitutes an alphabet, such as being an ordered sequence, does not appear to flexibly map to less familiar domains in LLMs like it does in people. Evidence for this comes from the Rule Check task, where we tested LLM performance on each of the rules separately. Here we see that the repetition rules could easily be applied to novel alphabets. This makes sense because if one were to create a function to repeat a character in a string this could be done without knowing the alphabet or the character. In contrast, the models had far more trouble with the predecessor and second successor rules, that, in order to solve correctly, both require that an alphabet is encoded as an ordered list of letters/symbols and an abstraction of what constitutes previous and next. This coincides with previous work, where Hodel and West, (2023) and Lewis and Mitchell, (2024) found that GPT models were only able to solve letter-string analogies when presented with familiar letters in their standard order. As soon as unfamiliar symbols were used or familiar letters were shuffled, performance dropped drastically. We noted that in the Greek domain the letters were also ordered, but in our Symbol domain they were not, which could perhaps explain why Greek items were easier. So, to check whether order was also a factor in our Symbol domain, we adapted the task to make the Symbol alphabet ordered by their unicode values. However, this adaptation resulted only in some improvement in the Claude and Gemma models, and our findings still held (see Appendix E).
@@
We also investigated which kinds of errors people and LLMs made. This is important because letter-string analogies, like many four-term visual analogies apply ambiguous rules (e.g., Opiełka et al.,, 2024), and can be solved correctly in multiple ways Hofstadter and Mitchell, (1994). The two main ways to solve the items in our task were what we considered the “correct” way (e.g., a b : a c c :: g h : g i i) and the “literal” way (e.g., a b : a c c :: g h : g c c). People did not use the “literal” rule, whereas the models all did to varying degrees (ranging from 5-21%). The other main difference between human and LLM errors, was that children’s erroneous responses were generally more distant (measured with Levenshtein string distance) from the “correct” response than those of LLMs. This could be because children reverted to associative strategies that we didn’t account for in our error coding scheme, given that this is the first time letter-string analogies have been administered to children. It will be interesting for future work to investigate children’s development and ‘shift’ from associative to relational responses on letter-string analogies, especially compared to verbal and visual domains Stevenson and Hickendorff, (2018); Gentner, (1988).
@@
Based on our results and previous work, these LLMs appear to have brittle, inflexible abstractions of what represents an alphabet in the context of the letter-string analogy task, despite being given the ordered list of letters/symbols before each item. In addition, LLMs may not have robust abstractions for predecessor and successorship Opiełka et al., (2025). It appears that LLMs like GPT-4 can only perform these abstractions by creating and executing code to map the novel alphabet to new positions Webb et al., (2024) and computing previous and next letters. This is of course very different from how children solve such problems.
@@
In contrast, in children, familiarity with letters or symbols does not seem to influence how well they solve letter-string analogies. As such, our results add to the accumulating evidence that questions whether reasoning actually occurs in these LLMs (Wu et al.,, 2024; Gendron et al.,, 2024). Interestingly, in 1980, Schank concluded that there wasn’t much intelligence in artificial intelligence given its limited ability to generalize. Similarly, Doumas et al. 2022 argue that robust analogical transfer is a uniquely human ability. Based on our findings so far we concur, and now ask the question: Is generalization to unfamiliar domains indeed what separates human general intelligence from that of artificial general intelligence? The challenge now is to create uncontaminated far generalization tasks that AI models have not been trained on to answer this question.
Research Transparency Statement
Artificial intelligence
No artificial intelligence assisted technologies were used in this research or the creation of this article.
Funding Information
This information is excluded to allow for anonymous review.
Open Science Practices

### v3

@@ 3 Results
We use mixed ANOVAs to (1) compare performance between our between-subjects participant groups (Adults, Children, and each of the LLMs) on the Latin alphabet and (2) test whether each participant group could generalize analogy solving by performing similarly across alphabets (i.e., our repeated within-subjects factor). All plots show the means and standard errors as error bars.
@@
We expected LLMs to be able to solve letter-string analogies with the Latin alphabet at the same level as adults (Webb et al., 2023) and that both adults and LLMs would outperform children (Thibaut and French, 2016) (hypotheses H1a-c). Similar to what we expected, adults and some LLMs, except Google’s Gemma-2 27B and Anthropic’s Claude 3.5, performed better than children in the Latin alphabet domain. Open AI’s GPT-4o performed similarly to adults, followed closely by Meta’s Llama-3.1 405B. See Figure 1 and Table 2 for more detailed results.
@@
Table 2: Descriptive statistics on letter-string analogy performance by Participant Group and Alphabet.
3.2 RQ2: How well do adults, children and LLMs generalize letter-string analogy solving from Latin to Greek (near) and Symbol (far) alphabets?
As expected, adults and children performed similarly across alphabets (see Figure 1). But, as we suspected, LLM performance indeed degraded in less familiar alphabets (ANOVA results shown in Table 3). More specifically, for each model, performance degraded significantly from the Latin to Greek alphabet (posthoc Bonferonni-corrected t-test results all p<.001, except for Llama-3.1 405Bp = 0.012) and then again from the Greek alphabet to the Symbol list (posthoc Bonferonni-corrected t-test results all p<.001).
Table 3: Post hoc ANOVA Results for main Alphabet effect by Participant Group
@@
Adults 1.59 96.9 0.95 1.000
Children 2.00 76.0 0.27 1.000
Claude-3.5 1.65 87.6 29.5 <.001
Gemma-2 27B 2.00 106.0 88.2 <.001
GPT-4o 2.00 100.0 55.0 <.001
Llama-3.1 405B 1.70 90.1 135.0 <.001
@@ 3.3 RQ3: Why can’t LLMs generalize letter-string analogy solving like children?
3.3.1 Performance by Item
To understand why the LLM’s had trouble generalizing letter-string analogy solving we examined item-by-item performance. Table 4 shows an overview. Here we see that the LLMs and humans perform best on item 1, that involves only the first successor transformation, and worst on item 5, that involves both the predecessor transformation and repetition generalization. Item 2 also involves the same repetition rule as item 5, but was solved better by LLMs and children; therefore, it appears that the predecessor rule is what gives both LLMs and children the most trouble. The other item people and LLMs have relatively more trouble with is item 3. This item involves the second successor rule. In sum, the predecessor and second successor rules appear to be the most difficult rules from our item set for people and LLMs to apply.
Table 4: Mean proportion correct (SD) by Participant Group for each Item
@@ 3.3.1 Performance by Item
3.3.2 Next-Previous Letter Task
We designed the Next-Previous Letter Task to check that the LLMs had the requisite knowledge of predecessor and successor to solve letter-string analogies. For this new task we provided an ordered list of letters/symbols and asked the LLMs what the previous and next two letters were given a specific letter. Each rule was tested 5 times, resulting in 20 items total.
We used this optimized prompt: Here is an ordered list of letters or symbols [Latin alphabet|Greek alphabet|Symbol list]. Which letter or symbol is [one|two] [before|after] [X]?. See Table 5 for rules and example items.
Table 5: Next-Previous Letter Task: Example Items From the Latin Alphabet.
X correct Rule
c d next_1
c e next_2
d c prev_1
e c prev_2
Figure 3: LLM performance by rule type across alphabet domains.
As can be seen in Figure 3, all models do best when asked to identify the next or previous letter and worse when it concerns identifying two before or two after. Furthermore, Claude-3.5 performed well and similarly in all three domains, which is in contrast to its letter-string analogy performance that degrades from baseline to near to far domains. Similarly, GPT-4o performs well on the Next-Previous task in the Latin and Greek domain, but in the analogy task, its performance degrades from Latin to Greek. For Llama-3.1 405B transfer from the Latin to Greek to Symbol domain is similar across tasks, where in both tasks it does well with the Latin and Greek alphabets, but not the Symbol alphabet. Gemma-2 27B’s performance is surprisingly more spotty here in the Greek domain than the Symbol domain. In conclusion, these results could explain why the LLMs have trouble with item 3, involving the second successor, but the results do not explain why they have trouble with item 5 involving the predecessor.
@@ 3.3.3 Rule Check Task
To better pinpoint why the LLMs had difficulty generalizing to other alphabets, we created a simplified version of the original analogy task that explicitly tested each rule in isolation. LLM prompts were exactly the same as in our original letter-string task, see Section 2.1.3.
The rules were: (1) successor_1, the next letter; (2) successor_2, letter two places after; (3) predecessor_1, the previous letter; (4) predecessor_2, letter two places before; (5) repetition_1, repeating the last letter and (6) repetition_2, repeating both letters. Each rule was tested five times. See Table 6 for examples.
Table 6: Rule Check Task: Example Items From the Latin Alphabet.
@@ A B C D Rule AB
As Figure 4 shows, the LLMs we tested can solve all rules in the Latin alphabet and have no problem with repetition rules in the Greek and Symbol domains. The successor and predecessor rules were solved to differing degrees in the Greek alphabet, with Claude-3.5 performing best followed by GPT-4o. All models had trouble with the successor_2 and predecessor rules in the Symbol alphabet, where only the successor_1 rule sometimes formed an exception. This makes sense given the predict-the-next-token goal that LLMs are trained on (McCoy et al., 2024).
Figure 4: Rule-specific performance across alphabet domains for LLMs.
@@ 3.3.4 Error Analysis
In general, when solving letter-string analogies there are often multiple rules that could underlie the change from A to B Hofstadter and Mitchell (1994). Because we use very short strings, there are generally only two clearly correct responses. We considered the rules that people would generally prefer when responding, to be “correct”, such as if ab changes to ac, then gh changes to gi. However, the literal rule of replacing the last letter with c, with response gc could also be considered correct.
@@
To examine errors in more detail we created a set of categories based on those from (Lewis and Mitchell, 2025) and extended these to account for common errors in children Stevenson and Hickendorff (2018). In the Literal rule category, the change from A to B is literally copied to C such as a b : a c c :: g h : g c c rather than providing the more common response of g i i. In the One rule category, the response is partially correct, but only (part of) one of the rules in the problem was applied, such as in responses to the previous example, g h h (only repetition applied) or g i (only successor applied). Partially correct responses are common in children when problem load supersedes processing capacity Stevenson and Hickendorff (2018). In the Incorrect rule category, one of the other rules from our item set (i.e., successor, predecessor, repetition) was applied; for example, if the successor rule was used instead of the predecessor rule. For the Copy rule, the A, B or C term was copied as copying the C-term is common in young children (Stevenson and Hickendorff, 2018; Opiełka et al., 2024). Finally, all remaining erroneous responses were placed in the Other rule category. Given that our task was less complex than in (Lewis and Mitchell, 2025) (i.e., shorter strings, fewer rules), we were able to automatically code these categories.
Table 7: Proportions of error categories by participant group. Note. 5% of children’s responses were empty.
@@
Table 7 shows that adults and children did not use the Literal rule, whereas all models used it sometimes. For Gemma-2 27B and GPT-4o the Literal rule was one of the most common error types. The One rule was used most often in errors by Claude-3.5. The Incorrect and Copy rules were not used very often by people or models. And the Other rules were used most often by all, except Claude-3.5.
@@
For each erroneous response we computed the Levenshtein string distance, also known as optimal string alignment distance, from the expected “correct” response to the given response. This distance counts the minimum number of edit operations (insertion, deletion, substitution) needed to change one string into the other. Here we investigate whether there are differences in mean Levenshtein distance between adults, children and LLMs for “erroneous” responses. Figure 5 shows that that the Levenshtein distance for “erroneous” responses is greater for children on all alphabets than for LLMs. For adults, this is only the case for the Symbol alphabet. For LLMs the Levenshtein distance hovers just under the 2 for all alphabets. Note also that the standard errors for LLMs are also much smaller, but this is because the adults and sometimes children (Greek, Symbol alphabets) had far fewer “erroneous” responses to sample from. These results tells us that when children provide “erroneous” answers their responses tended to differ largely from the expected response. For example, three children responded ‘m m’ to the item ‘If c d changes to b d, what does m m n n change to?’, which has a Levenshtein distance of 6 from the expected response ‘l l n n’. The LLMs tended to provide 1 or 2 expected letters and 1 or 2 unexpected ones. For example, on the same item (and its variants) six GPT-4o runs provided ‘l m m n’ as a response, with a Levenshtein distance of 2 from the expected response.
Figure 5: Mean Levenshtein string distance between incorrect and expected responses.
4 RQ4: What is the effect of model size on letter-string analogy performance?
In general, larger LLMs perform better on reasoning tasks than smaller LLMs (Wei et al., 2022; Huang and Chang, 2022). As Figure 6 shows, typical scaling laws generally appear to hold for how well LLMs generalize analogy solving in the letter-string domain. Especially in the Symbol domain, we observe a marked performance increase from smaller to larger models.
(a) Claude
(b) Gemma
(c) GPT
(d) Llama
Figure 6: Effect of LLM size on proportion correct in letter-string analogy solving across alphabets Latin, Greek and Symbol.
@@ 5 Discussion
Our main finding is that the LLMs we tested, using the same prompts given to children, were not able to generalize letter-string analogy solving like children can. LLMs perform at or above the level of children on letter-string analogies in the familiar Latin alphabet, but their performance on these same problems reduces somewhat when using the Greek alphabet (near transfer) and deteriorates greatly when using our Symbol alphabet (far transfer).
@@
Why can’t LLMs generalize when solving letter-string analogies? For some LLMs, this appears to be because they do not meet underlying requisites, such as indicating the predecessor or second successor of a letter in sequence. This would make sense given the predict-the-next-token goal that LLMs are trained on McCoy et al. (2024). We tested this using the Next-Previous letter task, where models were explicitly prompted with an ordered list of letters or symbols and asked to identify the (second) successor or predecessor to a given letter or symbol. These results explain why the LLMs have trouble with analogies involving a second successor. But, the LLMs had little trouble identifying predecessors in the Next-Previous Letter task, so these results do not fully explain why LLM performance degrades from Latin to Greek and Symbol domains.
@@
The problem with LLMs transfer from the Latin to other domains seems to lie in that the conceptual abstraction of what constitutes an alphabet, such as being an ordered sequence, does not flexibly map to less familiar domains like it does in people. Evidence for this comes from the Rule Check task, where we tested LLMs on each rule in isolation. Here repetition rules could easily be applied to novel alphabets. This makes sense because repeating a character in a string can be done without knowing the alphabet. In contrast, LLMs had more trouble with predecessor and second successor rules. Both require an alphabet that is encoded as an ordered list of letters/symbols and an abstraction of what constitutes previous and next. This result aligns with previous work where GPT models could solve letter-string analogies with familiar alphabets in their standard order, but for shuffled alphabets performance dropped drastically, whereas for people performance remained the same Hodel and West (2023); Lewis and Mitchell (2025). We noted that in the Greek domain the letters were also ordered by unicode value, but in our Symbol domain they were not, which could perhaps explain why Greek items were easier. So, to check whether order was also a factor in our Symbol domain, we adapted the task to make the Symbol alphabet also ordered by their unicode values. However, this adaptation resulted only in some improvement in the Claude and Gemma models, and our findings still held (see Appendix D).
@@
We also investigated which kinds of errors people and LLMs made. This is important because letter-string analogies, like many four-term visual analogies, apply ambiguous rules (e.g., Opiełka et al., 2024), and can be solved correctly in multiple ways Hofstadter and Mitchell (1994). The two main ways to solve the items in our task were what we considered the “correct” way (e.g., a b : a c c :: g h : g i i) and the “literal” way (e.g., a b : a c c :: g h : g c c). People did not use the “literal” rule, whereas the models all did to varying degrees (ranging from 5-21%). The other main difference between human and LLM errors, was that children’s erroneous responses were generally more distant (Levenshtein string distance) from the “correct” response than those of LLMs. This could be because children reverted to non-analogical strategies that we didn’t account for in our error coding scheme, given that this is the first time letter-string analogies have been administered to children.
@@
Based on our results, LLMs appear unable to create on-the-fly representations of novel alphabets in the context of the letter-string analogies as well as the next-previous letter task —despite being given the ordered list of letters/symbols before each item. This inability was clear for both larger and smaller LLMs, although relative performance did scale with model size. A possible explanation lies in work studying the internal representations of LLMs, where abstract concepts like "antonym" show invariant linear representations, but "previous" and "next" do not Opiełka et al. (2025). It appears that LLMs require an in-weight linear representation of an alphabet to successfully solve letter-string analogies. For novel alphabets, next-token-prediction does help them solve analogies with simple repetition and successor rules, but not with more complex rules and not at the level of children. Indeed, Webb et al. 2024 found that GPT-4 can only perform these abstractions by creating and executing code to map the novel alphabet to new positions and compute previous and next letters. This is of course very different from how children solve such problems.
@@
In contrast, our results show that in children, familiarity with letters or symbols does not influence letter-string analogy solving. As such, our results add to the accumulating evidence that questions whether reasoning actually occurs in these LLMs (Wu et al., 2024; Gendron et al., 2024; Razeghi et al., 2022). Interestingly, in 1980, Schank concluded that there wasn’t much intelligence in artificial intelligence given its limited ability to generalize. Similarly, Doumas et al. (2022) argue that robust analogical transfer is a uniquely human ability. Based on our findings so far we concur, and now ask the question: Is generalization to unfamiliar domains indeed what separates human general intelligence from that of artificial general intelligence? The challenge now is to create uncontaminated far generalization tasks that AI models have not been trained on to answer this question.
6 Acknowledgments
This research was funded by the the Dutch Research Council (NWO) project "Learning to solve analogies: Why do children excel where AI models fail?" with project number 406.22.GO.029 awarded to Claire Stevenson. We thank Veerle Vijverberg and Talea Sibum for their assistance with data collection. We also thank the University of Amsterdam CreAI Lab for fruitful discussions related to this work and the anonymous reviewers and the Action Editor for their helpful comments.
Appendix A Open Science Practices